OpenStudy (anonymous):

a^2+1^2=j^2, sovle for a

OpenStudy (anonymous):

(j-a)(j+a) = 1

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

If you are to solve for a then you would have a^2 = j^2 - 1 a = +-sqrt(j^2 - 1)